# Routing in Quantum Networks with End-to-End Knowledge

### v1

@@ Abstract
Given the diverse array of physical systems available for quantum computing and the absence of a well-defined quantum internet protocol stack, the design and optimisation of quantum networking protocols remain largely unexplored. To address this, we introduce an approach that facilitates the establishment of paths capable of delivering end-to-end fidelity above a specified threshold, without requiring detailed knowledge of the quantum network’s properties, which we call the “grey box approach”. In this study, we define algorithms that are specific instances of this approach and evaluate them in comparison to Dijkstra’s shortest path algorithm and a fully knowledge-aware algorithm through simulations. Our results demonstrate that one of the grey box algorithms consistently outperforms the other methods in delivering paths above the fidelity threshold, across various network topologies and the number of source-destination pairs involved, while maintaining significant levels of fairness among the users and being robust to inaccurate estimations of the expected end-to-end fidelity.
@@ I Introduction
  • We demonstrate that the proposed $kx_{0}$ algorithm (III-B) consistently outperforms other methods in achieving higher end-to-end fidelity across various network topologies and source-destination pairs employed while maintaining significant levels of fairness among users.
@@ II State of the art
Shi et al. [44] introduced the QCAST framework, which adopts a time-slot structure, a modified version of the approach proposed by Pant et al. [38]. Each time-slot consists of four phases. During phase one, all nodes receive information regarding the current source-destination pair intending to establish end-to-end entanglement. In phase two, routes are determined for the requests from phase one using the QCAST algorithm, and external entanglement links are established between neighbouring nodes, utilising the qubits and channels associated with these nodes. Phase three involves nodes exchanging link state information through classical channels. Finally, in phase four, each node establishes internal links to complete the route between the source and destination. That work has inspired us to define the system, illustrated in Section III, which also involves multiple phases in a time-slotted system. Zhang et al. [45] introduced a novel fragmentation-aware algorithm designed to address resource contention issues across multiple paths. The core concept involves redefining network resources as ’public’ (shared among all paths) rather than ’private’ (dedicated to specific paths), contrasting with the approach taken by QCAST. Zhao et al. [46] introduced a method aimed at enhancing network throughput over QCAST by as much as 68.75%, achieved through the provision of redundant external links established during phase two. Such optimisations are orthogonal to the problem addressed in this work and, in principle, they can be used in combination with the solutions proposed.
Despite its importance for distributed quantum computing applications, fidelity is often overlooked in the literature as a routing metric. Li et al. [43] proposed an iterative routing algorithm based on entanglement purification to provide fidelity-guaranteed paths. In [56] we proposed a vertex-weighted routing algorithm that delivers paths above certain threshold fidelity for heterogeneous nodes with mixed efficiency figures. Other routing metrics include latency [40, 34], fairness [41], and number of repeaters [42]. A common trait of above all routing metrics is that they require detailed knowledge of the network characteristics with certainty such as link quality or length [36], EPR generation rate [37], EPR success rate [44, 45, 46], purification success rate [43], and node quality [56]. In this work, we relax such an assumption and design a solution for path selection that only uses knowledge of the network topology, i.e., how the quantum repeaters are interconnected via quantum links, and an estimate of the end-to-end fidelity between end nodes.
@@ III Network and Motivation
In this section, we establish the network design and motivation for this work. In Section III-A, we provide a comprehensive discussion of the model of the system. In Section III-B, we establish the problem of routing in quantum networks.
@@ III-A System Model
We are aware that the number of possible paths in a generic network can be very large, which would render this approach practically unfeasible, but we make the following two observations to advocate such a solution, at least in some scenarios. First, the topology of a network is expected to remain stable for a time that is much longer than the time scales relevant for execution of distributed quantum computing applications, as modifying the topology implies the installation of a device (quantum computer or repeater) and its physical interconnection with the quantum network or provisioning additional infrastructures (satellite or optical fibre). Therefore, the information about end-to-end fidelity may be accumulated in the Storage over time through a background process, prioritising shorter paths between pairs of end nodes, which consume fewer resources and are more likely to exhibit a higher fidelity. Second, in a geographical quantum network, the number of quantum devices may be substantially smaller than the number of quantum repeaters that are needed to interconnect them, which would lead to a relatively small number of combinations for source/destination pairs, despite such a number being quadratic on the cardinality, i.e., $\mathcal{O}(N^{2})$, with $N$ quantum devices.
@@
where each pair of qubits in a link is assumed to be in a Werner state with initial fidelity $F$ upon generation of the local entanglement.
@@
At the time of composing this paper, Liu et al. [53] have recently published research that partially overlaps with the themes of our study. They introduced an approach for inter-domain routing among Autonomous Systems (ASs), utilising a methodology that involves selecting the best $k$ paths based on a randomised benchmarking procedure. This procedure estimates the end-to-end quality of each path [35]. Despite the apparent similarities, there are critical distinctions to be underscored. Primarily, our research focuses on intra-domain quantum routing, thus addressing a fundamentally different aspect of network management compared to the inter-domain routing proposed by Liu et al. [53]. Although our approach also prioritises the selection of the best $k$ paths, our criteria hinge on the underlying network topology, which is integral to solving intra-domain routing challenges. Moreover, we advance a more granular analysis of path fidelity by beginning with the fidelity assessments of depolarising channels at individual hops. Distinctively, our routing policies prioritise paths based on the minimum required fidelity rather than maximum fidelity. We demonstrate that this approach enhances resource utilisation significantly, setting our findings apart from those of Liu et al. [53].
@@
Typically, it is assumed that during the period the system will evolve as a sequence of evenly-spaced time slots, each including the phases from 3 to 6 in Figure 3: establish local entanglement in a link (3), perform Bell state measurements (4) transfer the classical results to one of the quantum devices so that it can perform $\sigma_{z}/\sigma_{x}$ corrections (5), and consume the qubits on the quantum devices by the distributed applications running on them (6). The duration of the time slot must be less than the decoherence time of the qubits, which with today’s state-of-the-art technology is very challenging. Furthermore, we note that local entanglements and Bell state measurements are stochastic processes, as they may fail with a given probability that depends on the specific technology used. During the operation of the system, these failures may happen in Phases 3 and 4 and affect the long-term throughput, i.e., the number of end-to-end entangled qubits that can actually be used by the source and destination devices in the unit of time. Instead, in the remainder of the paper, we will focus only on Phase 2, i.e., the route selection process, hence we can consider the stochastic failures orthogonal to our work.
@@ III-B The Routing Problem
We address the problem under the general framework illustrated by Algorithm 1, by proposing several specific policies as follows:
@@
Input: Network as a graph $G(V,E)$, E2E fidelity estimations, source-destination pairs $\{s_{k},d_{k}\},k=1..n_{sd}$, where $n_{sd}$ is the number of connections, $F_{th}$ is the fidelity threshold, and $\mathcal{A}$ is the path selection policy
@@
This policy makes use of additional knowledge of quantum repeaters i.e., their efficiency figures and is a white box approach solution to the quantum routing problem. Each node $v$ in the quantum network is assigned a cost $c(v)$ according to its efficiency figure. Then a vertex-weighted Dijkstra’s algorithm is employed for path-selection. A preliminary study of such a policy was done in [56]. Path selected:
@@
This policy selects the path with minimum end-to-end fidelity (yet above fidelity threshold) out of first $k$ shortest-paths between source and destination list, where $k$ is a configuration parameter of the algorithm. Note that those paths may be of different lengths (not all equal to the shortest one). The intuition behind this policy is that selecting paths with a lower fidelity, corresponding to quantum repeaters in low-quality classes, in the early rounds of the loop in Algorithm 1 should leave more resources available to connections under evaluation later on. Since no per-node/per-link information is needed, the $k$ shortest-path is a grey-box approach solution to the quantum routing problem. Path selected:
@@ IV Simulation Methodology
We developed a specialised Python-based simulator to compare the performance of the algorithms described in Section III-B, when applied to a network as described in Section III-A. Specifically, we aim at characterising the fidelity achieved by a target set of source-destination pairs, depending on the different routing strategies, when applied to phase 2 of Figure 3. As will be clear from the experimental results, the order in which paths are computed for the different source-destination pairs in the set may determine the achieved fidelity for a specific source-destination pair. Therefore, we use Monte Carlo simulations to generate random orders in which we compute the best paths for a given set of source-destination pairs, and average the results to remove the dependence on the position in the order.
With respect to the generic network shown in Figure 1, in the following we focus on two specific topologies, referred to as regular and random, respectively. In the regular topology, the quantum repeaters are arranged in a $n\times n$ square grid with a connection between top and bottom transport nodes to avoid any edge effects. The source and destination nodes are located on the opposite edge of the transport network grid. In the random topology, we use the well-known Waxman model [57] to generate the specific network topology. Specifically, we use the Waxman model to generate the network connecting the quantum repeaters among them. Then, we select $n_{sd}$ repeaters using a uniform random distribution and attach them to the source nodes, and another $n_{sd}$ repeaters to attach the destination nodes. When generating the network among repeaters according to. the Waxman model, a link between any two nodes $u$ and $v$ is added with the following probability:
Figure 4: Blocking probability(BP) vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
@@
Finally, a parameter $\theta$, path establishment order, is defined that represents the order (from 1 to the number of SD pairs considered in the experiments, $n_{sd}$) of a specific SD pair in the sequence used to serve all pairs.
Figure 5: Probability mass function (PMF) of path length with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.
IV-A Performance figures
@@
First, we consider the Blocking Probability (BP) across the entire set of considered SD pairs, computed as the fraction of SD pairs to which the routing algorithm assigns a path, over the total number of SD pairs requesting connectivity. In addition, for some specific analyses, we also consider the Blocking Probability per edge (BP/E) defined as the Blocking Probability divided by the number of edges in the network.
A second index we use is the fidelity achieved by SD pairs over the paths allocated by the routing algorithm, computed as in Equation 3.
@@ V Results
In this section, we evaluate the performance of the system presented in Section III, and specifically the performance of different routing approaches presented in Section III-B. Throughout this study, a quantum network consisting of 25 quantum repeaters is used irrespective of topology and number of source-destination pairs involved. A total of 100 random combinations of source and destination nodes are considered in addition to 100 random assignations of HQ and LQ nodes to the transport network. This leads to a total of $10,000\times n_{sd}$ potential paths i.e., $50,000$ for $n_{sd}=5$ for each $\xi$ value. The performance results for each $\xi$ are obtained by averaging over random assignation of HQ and LQ along with t-Student’s 95% confidence intervals. The results will only be shown for $10.5<\xi<$ 1 as having quantum network with $\xi<0.5$ only further decreases network performances since it would involve more lower quality nodes than higher quality nodes. The fidelity of the initially shared EPR pairs should be as high as possible, ideally close to 1, to tolerate the reduction in fidelity caused by noisy entanglement swapping procedures, thus ensuring that the entanglement shared between the source-destination pair remains above the fidelity threshold. However, achieving a fidelity of 1 is unrealistic; therefore, we assume the initial EPR pairs shared among neighbouring nodes have a fidelity of 0.975. We set the efficiency figures for high-quality and low-quality nodes at $\eta_{h}=0.999$ and $\eta_{l}=0.8$, respectively. It is important to note that the typical fidelity required by quantum applications should exceed 0.5. Consequently, we establish the threshold fidelity required for quantum applications at 0.53. Additionally, we set $k=10$ for both the $k$-shortest path and $kx$ path-selection approaches, providing a sufficient number of shortest paths from which to select. The results are categorised into six categories i.e., blocking probability, fairness, number of source-destination pairs, average degree, robustness, and range of efficiency figures. For each of these metrics, we evaluate the performance of all the path selection policies defined in Section III-B: shortest-path (SP) and knowledge-aware (KA), as baselines, and $k$-shortest path (KSP) and $kx$-path selection, as the proposed novel approaches.
@@
Figure 6: Blocking probability(BP) vs. Fraction of high-quality nodes ($\xi$) in the transport network with a fidelity threshold ($F_{th}$) = 0.53, and number of source and destination ($n_{sd}$) = 5.

### v2

@@ Abstract
Given the diverse array of physical systems available for quantum computing and the absence of a well-defined quantum internet protocol stack, the design and optimisation of quantum networking protocols remain largely unexplored. To address this, we introduce an approach that facilitates the establishment of paths capable of delivering end-to-end fidelity above a specified threshold, without requiring detailed knowledge of the quantum network’s properties. In this study, we define algorithms that are specific instances of this approach and evaluate them in comparison to Dijkstra’s shortest path algorithm and a fully knowledge-aware algorithm through simulations. Our results demonstrate that one of the proposed algorithms consistently outperforms the other methods in delivering paths above the fidelity threshold, across various network topologies and the number of source-destination pairs involved, while maintaining significant levels of fairness among the users and being robust to inaccurate estimations of the expected end-to-end fidelity.
@@ I Introduction
  • We demonstrate that the proposed $kx_{0}$ algorithm (III-E) consistently outperforms other methods in achieving higher end-to-end fidelity across various network topologies and source-destination pairs employed while maintaining significant levels of fairness among users.
@@ II State of the art
Shi et al. [44] introduced the QCAST framework, which adopts a time-slot structure, a modified version of the approach proposed by Pant et al. [38]. Each time-slot consists of four phases. During phase one, all nodes receive information regarding the current source-destination pair intending to establish end-to-end entanglement. In phase two, routes are determined for the requests from phase one using the QCAST algorithm, and external entanglement links are established between neighbouring nodes, utilising the qubits and channels associated with these nodes. Phase three involves nodes exchanging link state information through classical channels. Finally, in phase four, each node establishes internal links to complete the route between the source and destination. That work has inspired us to define the system, illustrated in Section III, which also involves multiple phases in a time-slotted system. Zhang et al. [45] introduced a novel fragmentation-aware algorithm designed to address resource contention issues across multiple paths. The core concept involves redefining network resources as ‘public’ (shared among all paths) rather than ‘private’ (dedicated to specific paths), contrasting with the approach taken by QCAST. Zhao et al. [46] introduced a method aimed at enhancing network throughput over QCAST by as much as 68.75%, achieved through the provision of redundant external links established during phase two. Such optimisations are orthogonal to the problem addressed in this work and, in principle, they can be used in combination with the solutions proposed.
Despite its importance for distributed quantum computing applications, fidelity is often overlooked in the literature as a routing metric. Li et al. [43] proposed an iterative routing algorithm based on entanglement purification to provide fidelity-guaranteed paths. In [56] we proposed a vertex-weighted routing algorithm that delivers paths above certain threshold fidelity for heterogeneous nodes with mixed efficiency figures. Other routing metrics include latency [40, 34], fairness [41], and number of repeaters [42].
A common trait of above all routing metrics is that they require detailed knowledge of the network characteristics with certainty such as link quality or length [36], EPR generation rate [37], EPR success rate [44, 45, 46], purification success rate [43], and node quality [56]. We classify approaches that construct routing metrics based on detailed network information as “white-box” approaches. In this work, we relax such an assumption and design a solution for path selection that only uses knowledge of the network topology, i.e., how the quantum repeaters are interconnected via quantum links, and an estimate of the end-to-end fidelity between end nodes. We term this method a “grey-box” approach, as it does not demand precise network information but relies on select, relevant insights.
@@ III Network and Motivation
In this section, we establish the network design and motivation for this work. In Section III-A, we provide a comprehensive discussion of the model of the system. In Section III-B, we discuss end-to-end fidelity estimation in the grey-box approach. In Section III-C, we discuss end-to-end entanglement. In Section III-D, we discuss the slotted model of the time evolution. In Section III-E, we establish the problem of routing in quantum networks.
@@ III-B End-to-end fidelity estimation
We are aware that the number of possible paths in a generic network can be very large, which would render this approach practically unfeasible, but we make the following observation to advocate such a solution, at least in some scenarios. The topology of a network is expected to remain stable for a time that is much longer than the time scales relevant for execution of distributed quantum computing applications, as modifying the topology implies the installation of a device (quantum computer or repeater) and its physical interconnection with the quantum network or provisioning additional infrastructures (satellite or optical fibre). Therefore, the information about end-to-end fidelity may be accumulated in the Storage over time through a background process, prioritising shorter paths between pairs of end nodes, which consume fewer resources and are more likely to exhibit a higher fidelity. However, it is crucial to consider that network topology information can be influenced by environmental factors beyond physical layout changes. Such environmental interactions may alter link performance, thereby impacting the accuracy of end-to-end fidelity predictions.
Consequently, developing efficient methods to manage the accumulation and updating of end-to-end fidelity knowledge remains an essential area for further research, as noted in recent studies (e.g., [35]). This topic, while outside the scope of this work, is complementary to our approach. In the meantime, given the early stages of development in this direction, we have implemented robustness tests to evaluate our methods under conditions with unreliable or outdated end-to-end knowledge, as will be discussed in subsequent sections.
@@ III-C End-to-end entanglement
where each pair of qubits in a link is assumed to be in a Werner state with initial fidelity $F$ upon generation of the local entanglement. It is worth noting that this equation models the heterogeneity across quantum repeaters, where each class of quantum repeater has a distinct efficiency figure $\eta_{g}$. The end-to-end fidelity $F_{p}$ thus becomes a function of both the number of repeaters in each class and their respective efficiency figures. This heterogeneity influences the cumulative fidelity of entanglement over the entire route, thereby impacting the performance of algorithms designed to optimize routing paths in quantum networks. In addition, the initial fidelity of EPR pairs in quantum links can vary, leading to heterogeneity in quantum links that impacts the end-to-end fidelity calculation. This end-to-end fidelity can be directly utilized in the grey-box algorithms to guide path selection and optimization. However, we assume uniform initial fidelity of EPR pairs $F$ across all quantum links in our analysis for simplicity.
@@
At the time of composing this paper, Liu et al. [53] have recently published research that partially overlaps with the themes of our study. They introduced an approach for inter-domain routing among Autonomous Systems (ASs), utilising a methodology that involves selecting the best $k$ paths based on a randomised benchmarking procedure. This procedure estimates the end-to-end quality of each path [35]. Despite the apparent similarities, there are critical distinctions to be underscored. Primarily, our research focuses on intra-domain quantum routing, thus addressing a fundamentally different aspect of network as compared to the inter-domain routing proposed by Liu et al. [53]. Although our approach also prioritises the selection of the best $k$ paths, our criteria hinge on the underlying network topology, which dictates the probability mass function (PMF) of path lengths and their corresponding end-to-end fidelities, which is integral to solving intra-domain routing challenges. Unlike the inter-domain framework proposed by Liu et al., which manages connectivity across multiple networks, our work is specifically designed for single-network, intra-domain scenarios where the routing scheme leverages detailed internal network knowledge rather than external network relationships. Moreover, we advance a more granular analysis of path fidelity by beginning with the fidelity assessments of depolarising channels at individual hops. Distinctively, our routing policies prioritise paths based on the minimum required fidelity rather than maximum fidelity. We demonstrate that this approach enhances resource utilisation significantly, setting our findings apart from those of Liu et al. [53].
III-D Slotted model
@@
Typically, it is assumed that during the period the system will evolve as a sequence of evenly-spaced time slots, each including the phases from 3 to 6 in Figure 3: establish local entanglement in a link (3), perform Bell state measurements (4), transfer the classical results to one of the quantum devices so that it can perform $\sigma_{z}/\sigma_{x}$ corrections (5), and consume the qubits on the quantum devices by the distributed applications running on them (6). The duration of the time slot must be less than the decoherence time of the qubits, which with today’s state-of-the-art technology is very challenging. Furthermore, we note that local entanglements and Bell state measurements are stochastic processes, as they may fail with a given probability that depends on the specific technology used. During the operation of the system, these failures may happen in Phases 3 and 4 and affect the long-term throughput, i.e., the number of end-to-end entangled qubits that can actually be used by the source and destination devices in the unit of time. Instead, in the remainder of the paper, we will focus only on Phase 2, i.e., the route selection process, hence we can consider the stochastic failures orthogonal to our work.
@@ III-E The Routing Problem
We address the problem under the general framework illustrated by Algorithm 1, by proposing several specific policies. The framework takes as input the quantum network topology represented as a graph $G(V,E)$, the end-to-end (E2E) fidelity estimations, and a set of source-destination pairs $\{s_{k},d_{k}\}$ (where $n_{sd}$ represents the total number of such pairs). Additionally, a path selection policy $\mathcal{A}$ is employed to compute the optimal paths for the connections.
The algorithm begins by initializing the set of assigned paths $P_{a}$ to an empty set (line 1). For each request to establish entanglement between source-destination pairs $\{s,d\}$ (line 2), the policy $\mathcal{A}$ is used to compute a feasible path $\pi$ between the source and destination nodes in the network graph $G(V,E)$, considering the network’s current state and fidelity constraints (line 3). If a path $\pi$ is successfully found, it is added to the set $P_{a}$ (line 4). If no such path can be established, the connection between the source-destination pair $\{s,d\}$ is marked as blocked for that period (line 5).
The process repeats for all $n_{sd}$ pairs, and at the end of the algorithm, the final set of assigned paths $P_{a}$ is returned as the output (line 7). The policies are as follows:
@@
Input: Network as a graph $G(V,E)$, E2E fidelity estimations, source-destination pairs $\{s_{k},d_{k}\},k=1..n_{sd}$, where $n_{sd}$ is the number of connections, and $\mathcal{A}$ is the path selection policy
@@
This policy uses additional knowledge of quantum repeaters i.e., their efficiency figures. Therefore, it is a white-box approach solution to the quantum routing problem. Each node $v$ in the quantum network is assigned a cost $c(v)$ according to its efficiency figure. Then a vertex-weighted Dijkstra’s algorithm is employed for path-selection. A preliminary study of such a policy was done in [56]. Path selected:
@@
This policy selects the path with minimum end-to-end fidelity (yet above fidelity threshold) out of first $k$ shortest-paths between source and destination list, where $k$ is a configuration parameter of the algorithm. Note that those paths may be of different lengths (not all equal to the shortest one). The intuition behind this policy is that selecting paths with a lower fidelity, corresponding to quantum repeaters in low-quality classes, in the early rounds of the loop in Algorithm 1 should leave more resources available to connections under evaluation later on. Since no per-node/per-link information is needed, the $k$ shortest-path is a grey-box approach solution to the quantum routing problem. A more descriptive name for this algorithm could be “min-k-shortest path”, as it clarifies that the algorithm does not involve random selection among shortest paths or selection of multiple paths. However, for simplicity, we retain the term “$k$-shortest path” throughout the paper. Path selected:
@@ IV Simulation Methodology
We developed a specialised Python-based simulator to compare the performance of the algorithms described in Section III-E, when applied to a network as described in Section III-A. Specifically, we aim at characterising the fidelity achieved by a target set of source-destination pairs, depending on the different routing strategies, when applied to phase 2 of Figure 3. As will be clear from the experimental results, the order in which paths are computed for the different source-destination pairs in the set may determine the achieved fidelity for a specific source-destination pair. Therefore, we use Monte Carlo simulations to generate random orders in which we compute the best paths for a given set of source-destination pairs, and average the results to remove the dependence on the position in the order.
(a) Random
(b) Regular
Figure 4: Topology of the quantum network for $n_{sd}=5$ and 25 grid nodes. Source nodes $s_{n}$ (in red), destination nodes $d_{n}$ (in light blue), and grid nodes (in dark blue). In a regular network structure, the source-destination (SD) pairs are positioned farther apart, leading to a probability mass function (PMF) with less diversity in path lengths and a higher proportion of longer paths compared to a random network.
IV-A Topology and devices
With respect to the generic network shown in Figure 1, in the following we focus on two specific topologies, referred to as regular and random, respectively as shown in Figure 4. In the regular topology, the quantum repeaters are arranged in a $n\times n$ square grid with a connection between top and bottom transport nodes to avoid any edge effects. The source and destination nodes are located on the opposite edge of the transport network grid. In the random topology, we use the well-known Waxman model [57] to generate the specific network topology. Specifically, we use the Waxman model to generate the network connecting the quantum repeaters among them. Then, we select $n_{sd}$ repeaters using a uniform random distribution and attach them to the source nodes, and another $n_{sd}$ repeaters to attach the destination nodes. When generating the network among repeaters according to the Waxman model, a link between any two nodes $u$ and $v$ is added with the following probability:
@@ IV-A Topology and devices
Finally, a parameter $\theta$, path establishment order, is defined that represents the order (from 1 to the number of source-destination (SD) pairs considered in the experiments, $n_{sd}$) of a specific SD pair in the sequence used to serve all pairs.
IV-B Performance figures
@@
First, we consider the Blocking Probability (BP) across the entire set of considered SD pairs, computed as the fraction of SD pairs to which the routing algorithm assigns a path, over the total number of SD pairs requesting connectivity. A lower blocking probability indicates better network performance. In addition, for some specific analyses, we also consider the Blocking Probability per edge (BP/E) defined as the Blocking Probability divided by the number of edges in the network. As with blocking probability, a lower BP/E reflects improved network performance. Additionally, we analyse the probability mass function (PMF) of path length, which describes the probability distribution of the path lengths selected to connect SD pairs. PMF varies with the routing approach, the network parameter $\xi$, and the network topology.
A second index we use is the fidelity achieved by SD pairs over the paths allocated by the routing algorithm, computed as in Equation 3. We also consider the path establishment order ($\theta$), which represents the sequence in which SD pairs are served. In an optimised approach, the path order should minimally impact network performance. However, as shown in the results, this behaviour may not hold, particularly when the number of source-destination (SD) pairs increases, which can reduce network performance.
@@ V Results
In this section, we evaluate the performance of the system presented in Section III, and specifically the performance of different routing approaches presented in Section III-E. Throughout this study, a quantum network consisting of 25 quantum repeaters is used irrespective of topology and number of source-destination pairs involved. A total of 100 random combinations of source and destination nodes are considered in addition to 100 random assignations of HQ and LQ nodes to the transport network. This leads to a total of $10,000\times n_{sd}$ potential paths i.e., $50,000$ for $n_{sd}=5$ for each $\xi$ value. The performance results for each $\xi$ are obtained by averaging over random assignation of HQ and LQ along with t-Student’s 95% confidence intervals. The results will only be shown for $10.5<\xi<$ 1 as having quantum network with $\xi<0.5$ only further decreases network performances since it would involve more lower quality nodes than higher quality nodes. The fidelity of the initially shared EPR pairs should be as high as possible, ideally close to 1, to tolerate the reduction in fidelity caused by noisy entanglement swapping procedures, thus ensuring that the entanglement shared between the source-destination pair remains above the fidelity threshold. However, achieving a fidelity of 1 is unrealistic; therefore, we assume the initial EPR pairs shared among neighbouring nodes have a fidelity of 0.975. We set the efficiency figures for high-quality and low-quality nodes at $\eta_{h}=0.999$ and $\eta_{l}=0.8$, respectively. It is important to note that the typical fidelity required by quantum applications should exceed 0.5. Consequently, we establish the threshold fidelity required for quantum applications at 0.53. However, applications running on end nodes may demand higher fidelity levels. Our choice of a lower fidelity threshold enables clearer differentiation of algorithmic performance and enhances the interpretability of the results. Notably, as observed in a prior study [56], algorithm performance may converge at higher fidelity thresholds. For readers interested in outcomes based on these higher thresholds, we refer them to our preliminary analysis where we explored this phenomenon. Additionally, we set $k=10$ for both the $k$-shortest path and $kx$ path-selection approaches, providing a sufficient number of shortest paths from which to select. The results are categorised into six categories i.e., blocking probability, fairness, number of source-destination pairs, average degree, robustness, and range of efficiency figures. For each of these metrics, we evaluate the performance of all the path selection policies defined in Section III-E: shortest-path (SP) and knowledge-aware (KA), as baselines, and $k$-shortest path (KSP) and $kx$-path selection (with $x=0$ as $kx_{0}$ and $x=1$ as $kx_{1}$), as the proposed novel approaches.
@@
Figure 5: Blocking probability (BP) as a function of the fraction of high-quality nodes ($\xi$) in a randomized transport network. Parameters are set with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.
Figure 6: Probability mass function (PMF) of path length in a randomized transport network with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.
Figure 7: Blocking probability (BP) as a function of the fraction of high-quality nodes ($\xi$) in a regular transport network. Parameters are set with a fidelity threshold ($F_{th}$) of 0.53 and a source-destination count ($n_{sd}$) of 5.